# On $\mathcal{F}$-semi-transitive families of operators

### v6

@@ 5.2. Applications
In sequel, we assume that $\Omega=\mathbb{R}.$ and $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [26], for $\tau\in C_{b}(\mathbb{R}),$ we put
@@
[26, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
@@
[26, Lemma 3.3] For $\tau\in C_{b}(\mathbb{R})$ and each $f\in\mathcal{C}_{\tau}$, there is a sequence $(g_{n})$ in the set $\{\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}:\,\,0<\epsilon_{2}<\epsilon_{1}<1,\,\,K_{\epsilon_{2}}^{(\tau)}\text{ is compact}\}$ such that $fg_{n}\rightarrow f$ in $\mathcal{C}_{\tau}$.
@@
[26, Lemma 3.5] For each $k\in\mathbb{Z}$, $f\circ\alpha\in\mathcal{C}_{\tau\circ\alpha^{k+1}}$ if and only if $f\in\mathcal{C}_{\tau\circ\alpha^{k}}$. Moreover, in this case, $\|f\circ\alpha\|_{\tau\circ\alpha^{k+1}}=\|f\|_{\tau\circ\alpha^{k}}$.
@@
$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$
@@
Example 5.10.
Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 5.5 and Corollary 5.9 are satisfied in this case.

### v7

@@ 3.1. Applications to generalized weighted shifts on Hilbert C*-modules
Example 3.4.
Let $N=1,r_{1}=1$ and choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\alpha(t)=t-1$ for all $t\in\mathbb{R}.$ If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; or if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then the conditions of Corollary 3.3 are satisfied in this case.
3.2. Applications to generalized weighted composition operators on spaces of operator-valued continuous functions
In this subsection, we let $C_{0}(\Omega,\mathcal{C})$ denote the non-unital normed algebra of all continuous $\mathcal{C}$-valued functions on $\Omega$ vanishing at infinity and equipped with the norm
$\|f\|_{\infty}:=\sup_{x\in\Omega}\|f(x)\|$
and $C_{b}(\Omega,B(H))$ be the unital normed algebra consisting of all uniformly bounded $B(H)$-valued continuous functions on $\Omega$ equipped with the same norm $\|\cdot\|_{\infty}$.
Obviously, $C_{0}(\Omega,\mathcal{C})$ is (left) ideal in $C_{b}(\Omega,B(H))$ and the condition $(E)$ is satisfied. We will therefore denote $C_{0}(\Omega,\mathcal{C})$ and $C_{b}(\Omega,B(H))$ by $\mathcal{A}$ and $A_{1},$ respectively. For every triple $(K_{1},K_{2},m)$ with $K_{1}\subseteq K_{2}\subseteq\Omega,\,K_{1},K_{2}\text{ compact, }m\in\mathbb{N},$ let $\tilde{P}_{(K_{1},K_{2},m)}\in C_{0}(\Omega,\mathcal{C})$ be given by
$\tilde{P}_{(K_{1},K_{2},m)}(x)=u_{(K_{1},K_{2})}(x)P_{m}\quad\text{for all }x\in\Omega.$
We have the following lemma.
Lemma 3.5.
The set $\{\tilde{P}_{(K_{1},K_{2},m)}\}_{K_{1}\subseteq K_{2}\in\Omega,K_{1},K_{2}\text{ compact, }m\in\mathbb{N}}$ satisfies the condition $(\mathcal{P})$.
Proof.
Let $O$ be a non-empty open subset of $\mathcal{A}$ and $f\in O$. Then there exists some $\varepsilon>0$ such that $\varepsilon$– neighbourhood of $f$ is contained in $O$ as well. Since $f\in\mathcal{A},$ there exists some compact subset $K_{1}$ of $\Omega$ such that $\|f(x)\|<\frac{\varepsilon}{2}$ for all $x\in\Omega\setminus K_{1}$. Now, by continuity of $f,$ for each $x\in K_{1}$ we can find some open set $O_{x}$ of $\Omega$ containing $x$ such that $\|f(u)-f(x)\|<\frac{\varepsilon}{2}$ whenever $y\in O_{x}$. Then for all $m\in\mathbb{N}$ and each $y\in O_{x}$ we obtain
$\|(I-P_{m})f(y)\|\leq\|(I-P_{m})(f(y)-f(x))\|+\|(I-P_{m})f(x)\|$
$\leq\|f(y)-f(x)\|+\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}+\|(I-P_{m})f(x)\|$
By [21, Proposition 2.2.1], since $f(x)\in\mathcal{C},$ we can find sufficiently large $m_{x}\in\mathbb{N}$ such that $\|(I-P_{m})f(x)\|<\frac{\varepsilon}{2}$ for all $m\geq m_{x}.$ Hence, $\|(I-P_{m})f(y)\|<\varepsilon$ each $y\in O_{x}$ and all $m\geq m_{x}.$
The collection $\{O_{x}\}_{x\in K_{1}}$ is an open cover of $K_{1}$, hence there exist a finite subcover $\{O_{x_{1}},\ldots,O_{x_{M}}\}.$ Let $\tilde{m}:=\max\{m_{x_{1}},\ldots,m_{x_{M}}\}.$ Then, by the above we get
$\|\tilde{P}^{3}_{(K_{1},K_{2},m)}(x)f(x)-f(x)\|=\|(I-P_{m})f(x)\|<\varepsilon$ for all $x\in K_{1}$ and $m\geq\tilde{m}$. If $x\in\Omega\setminus K_{1},$ then for all $m\in\mathbb{N}$ we get
$\|\tilde{P}^{3}_{(K_{1},K_{2},m)}(x)f(x)\|+\|f(x)\|=\|u^{3}_{(K_{1},K_{2})}(x)P_{m}f(x)\|+\|f(x)\|\leq 2\|f(x)\|<2\frac{\varepsilon}{2}=\varepsilon.$
Thus,
$\|\tilde{P}^{3}_{(K_{1},K_{2},\tilde{m})}f-f\|_{\infty}<\varepsilon,\text{ so }\tilde{P}^{3}_{(K_{1},K_{2},m)}f\in O.$
Now we consider families $\{\alpha_{n,1}\}_{n\in\mathbb{N}}\dots,\{\alpha_{n,N}\}_{n\in\mathbb{N}}$ of homeomorphisms of $\Omega.$ We will assume that for every compact subset $K$ of $X$ and every $S\in\mathcal{F}$ there exists some $F\in\mathcal{F}$ with $F\subseteq S$ such that
$K\cap\alpha_{n,\ell}(K)=\varnothing\quad\text{and}\quad\alpha_{n,\ell}^{-1}(\alpha_{n,s}(K))\cap K=\varnothing$
for all $n\in F$ and each distinct $s,\ell\in\{1,\dots,N\}.$ Also, we consider families of unitary operators $\{U_{n,1}\}_{n\in\mathbb{N}}\dots,\{U_{n,N}\}_{n\in\mathbb{N}}$ on $H.$ For every $n\in\mathbb{N}$ and $l\in\{1,\ldots,N\}$ we let then
$\Phi_{n,l}:\mathcal{A}_{1}\rightarrow\mathcal{A}_{1}\text{ be given by }\Phi_{t,l}(f):=U_{n,l}^{*}(f\circ\alpha_{n,l})U_{n,l}$
for all $f\in\mathcal{A}.$ Obviously the system $\{\Phi_{n,1}\}_{n\in\mathbb{N}},\dots,\{\Phi_{n,N}\}_{n\in\mathbb{N}}$ is a system of isometric isomorphisms on $\mathcal{A}$ which is disjoint aperiodic with respect to $\{\tilde{P}_{(K_{1},K_{2},m)}\}$ and, clearly, the condition ($R$) is satisfied. Let $\{w_{n,1}\}_{n\in\mathbb{N}},\dots,\{w_{n,N}\}_{n\in\mathbb{N}}$ be families in $\mathcal{A}_{1}$ such that $w_{n,l}(x)$ is invertible for all $x\in\Omega$, $n\in\mathbb{N}$, $l\in\{1,\ldots,N\}$, and such that $\{w^{-1}_{n,1}\}_{n\in\mathbb{N}},\dots,\{w^{-1}_{n,N}\}_{n\in\mathbb{N}}$ also belong to $\mathcal{A}_{1}$ i.e.
$\sup_{x\in\Omega}\|w^{-1}_{n,l}(x)\|<\infty\quad\text{for all }n\in\mathbb{N}\text{ and }l\in\{1,\ldots,N\}.$
For each $n\in\mathbb{N}$, and $l\in\{1,\ldots,N\},$ let then $b_{n,l}:=w_{n,l}U_{n,l}$. Since each $b_{n,l}$ is invertible in $\mathcal{A}_{1}$, it follows that the condition $(C)$ is satisfied in this case.
Corollary 3.6.
Under the above notation and assumptions, if there exist dense subsets $H_{0},H_{1},\ldots,H_{N}$ of $H$ such that for each $s,l\in\{1,\ldots,N\}$ and every compact subset $K$ of $\Omega$ we have that
$\lim_{n\to\infty}\left[\left(\sup_{x\in K}\|w_{n,s}^{-1}(x)h_{s}\|\right)\cdot\left(\sup_{y\in K}\|(w_{n,\ell}\circ\alpha_{n,\ell}^{-1})(y)h_{0}\|\right)\right]=0$
for every $h_{s}\in H_{s}$ and $h_{0}\in H_{0},$ and, in addition, for each distinct $s,l\in\{1,\ldots,N\}$ and every $h_{s}\in H_{s}$ we have
$\lim_{n\to\infty}\left(\sup_{x\in K}\|(w_{n,l}\circ\alpha_{n,l}^{-1}\circ\alpha_{n,s})(x)\,w_{n,s}^{-1}(x)\,h_{s}\|\right)=0,$
then the families $\{T_{n,1}\}_{n\in\mathbb{N}},\ldots,\{T_{n,N}\}_{n\in\mathbb{N}}$ are $d\mathcal{F}$-semi-transitive.
Proof.
For each $s\in\{1,\ldots,N\}$, the set consisting of elements of the form
$u_{(K_{1},K_{2})}G_{s,m},$
where $G_{s,m}\in\mathcal{C}$ is such that
$G_{s,m}e_{i}\in H_{s}\text{ for all }j\in\{-m,\ldots,m\}\text{ and }G_{s,m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\ldots,m\}$
is obviously dense in $\mathcal{A}$ by Lemma 3.5 and the density of $H_{s}$ in $H$. (Here of course $m$ runs through the whole $\mathbb{N}$ and $(K_{1},K_{2})$ runs through all pairs of compact subsets $K_{1},K_{2}$ of $\Omega$ with $K_{1}\subseteq K_{2}.$) Let us denote this set by $\mathcal{A}_{s}$ for each $s\in\{1,\ldots,N\}$. Similarly, we let $\mathcal{A}_{0}$ be the set consisting of all elements of the form $u_{(K_{1},K_{2})}D_{m}$ where $D_{m}\in\mathcal{C}$ is such that
$D_{m}e_{j}\in H_{0}\text{ for all }j\in\{-m,\dots m\}\quad\text{and}\quad D_{m}e_{j}=0\text{ for }j\in\mathbb{Z}\setminus\{-m,\dots m\}.$
By some calculations similar to those in the proof of [16, Theorem 3.2] and [17, Proposition 2.7], it is then not hard to deduce that the conditions of Corollary 2.2 are satisfied. ∎
Example 3.7.
[…]
3.3. Applications to generalized weighted shifts on function sequence spaces
In this subsection, we assume that $\Omega=\mathbb{R}.$ and we let now $\mathcal{C}=C_{0}(\mathbb{R}).$ As in [15], for $\tau\in C_{b}(\mathbb{R}),$ we put
@@ 3.3. Applications to generalized weighted shifts on function sequence spaces
[15, Lemma 3.2] Let $\tau\in C_{b}(\mathbb{R})$ and $\epsilon_{1},\epsilon_{2}\in(0,1)$ with $\epsilon_{2}<\epsilon_{1}.$ Then for every compact subset $K_{\epsilon_{2}}^{(\tau)}\subseteq|\tau|^{-1}(\left[0,\epsilon_{2}\right]),$ there exists a function $\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}\in\mathcal{C}_{\tau}$ such that the followings hold:
@@
[15, Lemma 3.3] For $\tau\in C_{b}(\mathbb{R})$ and each $f\in\mathcal{C}_{\tau}$, there is a sequence $(g_{n})$ in the set $\{\mu_{K_{\epsilon_{2}}^{(\tau)},\epsilon_{1}}:\,\,0<\epsilon_{2}<\epsilon_{1}<1,\,\,K_{\epsilon_{2}}^{(\tau)}\text{ is compact}\}$ such that $fg_{n}\rightarrow f$ in $\mathcal{C}_{\tau}$.
@@
[15, Lemma 3.5] For each $k\in\mathbb{Z}$, $f\circ\alpha\in\mathcal{C}_{\tau\circ\alpha^{k+1}}$ if and only if $f\in\mathcal{C}_{\tau\circ\alpha^{k}}$. Moreover, in this case, $\|f\circ\alpha\|_{\tau\circ\alpha^{k+1}}=\|f\|_{\tau\circ\alpha^{k}}$.
@@
$\lim_{k\to\infty}\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\frac{1}{\prod_{i=1}^{r_{s}n_{k}}\big(w_{j+i-r_{s}n_{k}}^{(s)}\circ\alpha^{r_{s}n_{k}-i}\big)(t)}\Bigg)\cdot\Bigg(\sup_{t\in K_{\epsilon_{2}}^{(\tau\circ\alpha^{j})}}\prod_{i=1}^{r_{\ell}n_{k}}\big(w_{j+i}^{(\ell)}\circ\alpha^{-i}\big)(t)\Bigg)=0$